How many moles of gases are produced by the reaction of 6.65 kg of ammonium perchlorate? Report your answer to the nearest hundredth of a mole.
Answer #1

TThe decomposition reaction reaction of ammonium perchlorate can be written as

2NH4ClO4(s) -----> N2(g) + Cl2(g) + 2O2(g)+ 4H2O(l)

Number of moles of ammonium perchlorate= 6.65*000/117.5 = 56.6 moles.

From the equation 2 moles of ammonium perchlorate produces 4 moles of gases.

56.6 moles of ammonium perchlorate produces 56.6*4/2 = 113.19 moles of gases.